Factor
Don't understand what to do
5x^2+2x=0

Answer by stanbon(75887)

Factor
Don't understand what to do
5x^2+2x=0
---
Common factor: x
--
x(5x+2) = 0
---
Since the product is zero, one of the factors must be zero.
x = 0 or 5x+2 = 0
x = 0 or 5x=-2
x = 0 or x = -2/5
======================
cheers,
Stan H.